OpenStudy (anonymous):

Ramona has $8.00 in her savings account and adds $1.00 each week. Claire has $12.00 in her savings account and adds $3.00 each week. After how many weeks will Claire's account have twice as much money as Ramona's?

OpenStudy (anonymous):

Think of it as equations. Amount first put in + amount per week. So Ramona is 8 + 1x Claire is 12 + 3x You want to know at what point Claire will be 2(Ramona) so: 12+3x = 2(8+1x) and solve for x.

OpenStudy (djjo):

At the 4th week because: 8/12=9/15=10/18=11/21=12/24